LBC/LUCI Technical Observing, prep for Partner Science:
UT 2011 Apr 19

Observer: DThpmpson
Telescope GBechetti (19th) DGonzalez Huerta (20th):
Instrument Support: DOfficer (19th), JMorris (20th)
Telescope Support: MMidkiff (19th) JLittle (20th)
SW support:

Summary

19th lost to weather, high winds all night.

20th short delay in startup due to weather, then:
LBC
  • bias and darks
  • focal plane tilt
  • rotator center (no significant change)
  • guide/dither on SN09nr
  • Pointing co-pointing data (marred by problems with Blue)
LUCI
  • astrometric fields to check orientation
  • pointing data using pointing mask
  • rotcen on M92
  • science test on 1ES1927
  • darks

LUCI worked well, no problems with any mask motions. LBC had several blue controller errors that may make the pointing data useless. Convergence of dofpia was much quicker for blue than for red.

Details 20110419

01:10 Executing "WarmUp" on LBCs. LUCI mask exchange today. LBC hubs moved 1mm towards primaries. Some high clouds, 6.7C outside, M1 SX=4.9 DX=4.6 C. Windy, gusts over 25m/s. "Turn ON" of LBCs worked the first time. Connected to LBT.

Lost connection to internet, and thus wiki. Catch up from yesterday...

25Bias_Bino all files copied completely to archive
2Dark_Bino both files copied completely to archive

Switched to LUCI checkout. Move sieve mask in failed. Peter had to talk me through recovery because the internet connection was down. Ran flexure sequence with N3.75 and sent the plot to Walter.

Details 20110420

23:40 WarmUp and TurnON worked fine
Bias 00:07 25Bias_Bino run. Dome is not dark...one vent door is open to cool chamber, run MV. Counts are ~730 on Blue, 360 on Red. All transferred successfully to the archive. One thing though...with only 3 chips why are the red images the same size as the blue?

Dark 01:21 2Dark_Bino run. Two side vents open to equilibrate temps, so dome is not dark.

Clouds at sunset - no flatfields taken. Satellite loop looks like it will mostly clear off soon. Waiting a little at DGH's request.

Correct pointing and co-pointing ACT0184. B25559 R25556 before rangebal. B25906 R25903 after rangebal.

03:03 Collimate (dofpia, /first) Bl;ue converged quickly, red took 5-6 iterations. Clear all active optics etc because new global offsets were not implemented in PSF.

03:13 Correct pointing, co-pointing

03:15 DOFPIA, /first ... both converged in 4 cycles

Measure focal plane tilt, EL~51 for these observations (ACT0184)
  • 03:22 RB_rV_superfoc.ob
  • 03:27 RB_rV_superfoc180.ob
  • dofpia
  • 03:38 RB_rV_superfoc180.ob
  • 03:42 RB_rV_superfoc.ob

Repeat focal plane tilt measurements, EL~61 for these observations (ACT0222)
  • correct pointing, co-pointing at ACT0222
  • 03:52 dofpia
  • 03:57 RB_rV_superfoc.ob
  • 04:01 RB_rV_superfoc180.ob
  • dofpia
  • 04:12 RB_rV_superfoc180.ob
  • 04:16 RB_rV_superfoc.ob

Measure current rotcen values
  • Still at ACT0222
  • dofpia. Blue converged in 2, red took 6 iterations
  • rotatorcenter_pa0 B:43334 R:43201
  • rotatorcenter_pa180 B:43552 R:43546
  • dofpia
  • rotatorcenter_pa180 B:44223 R:044219
  • rotatorcenter_pa0 B:44449 R:44428

Red
Old rotator center: 1078. 2913.
New rotator center: 1078 2909
Xc=1086.3 Yc=2903.91 from geomap.
Xc=1076.89 Yc=2907.14 from averages.
Xc=1078.99 Yc=2909.4 from geomap.
Xc=1079.38 Yc=2909.23 from averages.

Blue
Old rotator center: 1035. 2924.
New rotator center: 1032 2924
Xc=1042.18 Yc=2919.69 from geomap.
Xc=1031.53 Yc=2922.99 from averages.
Xc=1023.38 Yc=2928.23 from geomap.
Xc=1032.47 Yc=2924.69 from averages.

  • No change to image headers for ROTCEN, differences small!
  • GEOMAP algorithm needs to be checked!

Guide and Dither
04:52 move to SN09 (13:10:58.90 11:29:29.60)
  • dofpia, /x2 (only 2-3 pupils found at defaults)
  • 05:00 ND_oldsn_sn09nr_BR.ob

Pointing Model Observations
  • Pointing log started
  • collimate
  • rangbal @ each position

Pointing Data:
ACT0365 53133 53130
WT10_343 54225 54222
ACT0397 55040 55207 AZ=-27
ACT0397 60433 60329 AZ=332 (unwrapped)
WT10_365 60852 60849 pos#1
WT10_408 61312 61309 pos#2
WT10_404 61701 61658
ACT0350 62307 62304
WT10_286 62752 62749 Red collim not good at low EL
ACT0203 63217 63213 Why is PA=-17 in co-pointing obs?
WT10_241 06:39 Blue CCD Controller error
199   2011/04/20   06:39:38.891284         OB execution error: none [src/lbc/actions.c:1145]
198   2011/04/20   06:39:38.773500         >>> preset/expose/save interrupted at preset time
197   2011/04/20   06:39:38.773432         camera system encountered a problem
196   2011/04/20   06:39:38.773379   CAMERA       CCD Controller error errcode:9 [src/camera/camera.c:1099]
195   2011/04/20   06:39:38.773325   CAMERA       reinitializing CCD Controller failed 

Turn off cameras and restart...came back OK, 14 minutes lost

And there is a hanging TFTP as before:
issuing >tftp -b32768 -o 192.168.0.201 put lbcb.20110420.060852.fits d:\lbcb.20110420.060852.fits > nul 2>&1< upload command failed with retcode:8

WT10_241 65020 65017
WT10_103 65433 65430
posn9 no source
WT10_245 07:02 Blue CCD Controller error
186   2011/04/20   07:02:08.653795         OB execution error: none [src/lbc/actions.c:1145]
185   2011/04/20   07:02:08.541246         >>> preset/expose/save interrupted at preset time
184   2011/04/20   07:02:08.541160         camera system encountered a problem
183   2011/04/20   07:02:08.541103   CAMERA       CCD Controller error errcode:9 [src/camera/camera.c:1099]
182   2011/04/20   07:02:08.541051   CAMERA       reinitializing CCD Controller failed 

Turn off cameras and restart...came back OK, Another 14 minutes lost

WT10_245 71310 71256
WT10_175 CCD Controller error
33   2011/04/20   07:15:47.839751         OB execution error: none [src/lbc/actions.c:1145]
32   2011/04/20   07:15:46.579597         >>> preset/expose/save interrupted at expose time
31   2011/04/20   07:15:46.203956         camera system encountered a problem
30   2011/04/20   07:15:46.203904   CAMERA       CCD Controller error errcode:9 [src/camera/camera.c:1185]
29   2011/04/20   07:15:46.203861   CAMERA       reinitializing CCD Controller failed
28   2011/04/20   07:15:46.203000   CAMERA    CCDCTRL    temperature safety limit exceded (4.0Celsius instead of -10.0Celsius)
27   2011/04/20   07:15:46.203000   CAMERA    CCDCTRL    detectors temperature: 277.2 [Kelvin] 

Turn off cameras and restart...nope...another error...
36   2011/04/20   07:21:37.314819   CAMERA       CCD Controller error, errcode:9 [src/camera/camera.c:556]
35   2011/04/20   07:21:37.313000   CAMERA    CCDCTRL    temperature safety limit exceded (61.3Celsius instead of -10.0Celsius)
34   2011/04/20   07:21:37.313000   CAMERA    CCDCTRL    detectors temperature: 334.4 [Kelvin] 

Turn off cameras and restart... again ...came back OK after auto-restart of HK, Another 15 minutes lost

WT10_175 73356 73254 This is pos#9
ACT0232 no star visible (pos#11) there is a cloud in that direction
ACT0317 We hit limits

We had many instances of error messages: "temperature safety limit exceded (-2.4Celsius instead of -10.0Celsius)" that seemed to trigger a hang or restart of the blue exposure. Each costs at least ~30s of lost time.

*Time's up for LBC. Because of the problems I am not at all confident that the last 2h15m are useful.

07:48 Switch to LUCI =================================

08:18 Collimating on BS9153. Take some image pairs on BS stars for astrometry.

All are Ks 5x10.0s N3.75
BS9153 0005 0006
BS9155 0007 0008
BS9160 0009 0010
BS9164 0011 0012 Seeing is blowing up!
BS9166 Failed to acquire guide star on several attempts, odd behavior, EL~74
BS9162 0013 0014 Seeing got better, cap cloud dissipated
BS9169 0015 0016
BS9177 0017 0018 Seeing highly variable.

09:07 Pointing model observations using pointing mask
Pointing log started, appending to the default log files.
Unwrap AZ so first point has AZ>360
Correct pointing

All are 3x2.0s to be able to see the reference circle
StarID LUCI# AZ cell
WT10_356 0023 388 01
ACT0405 0024 375 02
WT10_399 0025 338 03
ACT0341 0026 310 04
WT10_275 0027 280 05
WT10_190 0028 235 06
WT10_264 0029 208 07
WT10_077 0030 190 08
WT10_179 0031 140 09
WT10_236 0032 097 11
ACT0334 0033 065 12
WT10_332 0034 067 13
ACT0287 0035 087 14
ACT0253 0036 098 15
WT10_315 0038 093 16
ACT0213 0037 158 17
WT10_400 0039 012 02
WT10_399 0040 -22 03
WT10_334 0041 065 12
WT10_268 0042 123 10
ACT0094 0043 168 18
ACT0141 0044 186 19
WT10_298 0045 222 20
WT10_262 0046 247 21
WT10_224 0047 261 22
ACT0294 0048 277 23
WT10_343 0049 303 24
WT10_401 0050 349 25
WT10_404 0051 345 26
ACT0411 0052 366 27
WT10_380 0053 392 01
ACT0330 0054 388 28
WT10_371 0055 400 29
ACT0384 0056 393 30

Rotcen on M92 using pointing mask
  • Correct pointing, collimation nearby
  • 0058 at 0
  • 0059 at 180
  • 0060 at 270
  • 0061 at 90

1ES1927
  • Sci acq: 0062 sky, 0063 source, 0064 slit, 0065 src+slit. Offset -8.35, -2.56
  • J.spec 0066-0069, K.spec 0070-0073
  • Telluric acq: 0074-0076
  • Tell 0077-0078 J.spec 0079-0080 K.spec (lower EL than it should have been)
  • 0081-0083 lamps off K.flat
  • 0084-0086 halo1+2 K.flat
  • 0087-0089 lamps off J.flat
  • 0090-0092 halo1+2 J.flat

Exptime scaling on Sh2-106, various filters, for future PR image.

Darks 10x2.0s 5x10.0s 2x15.0s 1x120.0s 1x1800.0s

-- DavidThompson - 19 Apr 2011
Topic revision: r8 - 20 Apr 2011, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback